OpenStudy (kainui):

I wrote some code a while back to find the tangent of an angle, but now I'm not sure how it works, can someone help me figure it out again? haha

OpenStudy (kainui):

The mathematical formula is this: \[\Large \tan \theta = \frac{\sqrt{x^2+y^2}}{z}\] The reason I say it's code is because x, y, and z are all different values on the phone based on this picture http://www.larsivar.com/cp/images/android_AccForces.png and the angle is the angle between gravity and the normal vector to the face of the phone 

OpenStudy (kainui):

Here's a better picture of what I'm describing  So if you lay the phone straight down you have theta = 0, tangent =0 since z axis on the phone is parallel to the gravity vector (the only vector independent of the phone)